OpenStudy (anonymous):

Twenty-one more than five times a number is equal to the difference between 228 and four times the number. Find the number

OpenStudy (anonymous):

so 21 + 5*x = 228-4x 21 + 9x = 228 9x = 228-21 9x = 207 x = 23